### v1

# Expertized Caption Auto-Enhancement for Video-Text Retrieval

Abstract
The burgeoning field of video-text retrieval has witnessed significant advancements with the advent of deep learning. However, the challenge of matching text and video persists due to inadequate textual descriptions of videos. The substantial information gap between the two modalities hinders a comprehensive understanding of videos, resulting in ambiguous retrieval results. While rewriting methods based on large language models have been proposed to broaden text expressions, carefully crafted prompts are essential to ensure the reasonableness and completeness of the rewritten texts. This paper proposes an automatic caption enhancement method that enhances expression quality and mitigates empiricism in augmented captions through self-learning. Additionally, an expertized caption selection mechanism is designed and introduced to customize augmented captions for each video, facilitating video-text matching. Our method is entirely data-driven, which not only dispenses with heavy data collection and computation workload but also improves self-adaptability by circumventing lexicon dependence and introducing personalized matching. The superiority of our method is validated by state-of-the-art results on various benchmarks, specifically achieving Top-1 recall accuracy of 68.5% on MSR-VTT, 68.1% on MSVD, and 62.0% on DiDeMo.
1 Introduction
Video-text understanding has become a new mainstream research with the population of short videos on social networks. In virtue of the learning ability of large models, multi-modal pre-training models [11, 6] are developed and achieve initial success in video-text retrieval tasks. However, a stunning amount of data and computing resources are the essential basis to support the effectiveness of these pre-training models. Thus, it is argued that adapting the image-text pretrained model to video-text tasks is a lighter and more feasible approach [15, 31]. They adapt the visual embeddings to the video field by incorporating temporal information [21, 4] or refining the visual-text matching strategies [8, 31]. But unlike image-text pairs, video captions tend to be overly simplistic and one-sided, describing partial (even a signal) frames of the videos. These one-sided descriptions would be improperly matched to multifaceted video embeddings (Figure 1 (a)), resulting in biased cross-modal retrieval, and a bottleneck in performance improvement arises.
Figure 1: Comprehension gap between videos and texts
Frame sampling is one proper solution to handle the information imbalance issue. For instance, ClipBERT [10] sparsely samples video frames to meet the text descriptions. X-CLIP [16] further cuts text descriptions into words, employing cross-modal matching at both fine-grained (frame and word) and coarse-grained (video and sentence) levels. As video information is sparsely divided and broken, these methods are limited in comprehensively utilizing and exploring deep relations across videos and texts. Another branch to reduce the information gap is data agumentation. Extra data, such as HD-VILA-100M [30] and LF-VILA [21], are collected to expand the receptive field for model adaptation, thereby enriching matching knowledge of the two modalities. To eliminate the heavy data collection workload, later research [27] rewrites captions by lexical analysis, in which attributes are extracted and emphasized to enhance expressive ability in text modality. Large language models (LLMs), such as Chat-GPT and LLaMA [22], are additionally introduced for the rewriting tasks [26, 7]. Although up to 3% retrieval precision increase is achieved with the assistance of supplementary texts, the rewriting quality relies heavily on lexicon engineering [18] and hand-crafted prompts [17] (Figure 2 (a)). Recent works [2, 25] have been mired in the meticulous work of designing better prompts for caption augmentation, limited in human empiricism.
This paper proposes an Expertized Caption auto-Enhancement (ExCae) method, which adaptively generates and personalized select captions for video-text retrieval, releasing extra work of data acquisition and avoiding the trouble of manual prompt design. Specifically, a Caption Self-improvement (CSI) module is designed to derive captions with rich content and diverse perspectives from videos. A Prompt engineer is introduced in the CSI module, which gradually adjusts and improves the query prompts via self-learning. With the enriched expressions on the video-side (video+video-derived captions), cross-modal matching becomes a more challenging task. Thus, we further develop an Expertized Caption Selection (ECS) module consisting of multiple learnable experts to automatically specify appropriate expressions for cross-modal matching. Seeing Figure 1 (b), video representation space dilates with caption augmentation and the cross-modal information gap is reduced after expertized selection. This indicates that the proposed method improves the generalization ability through caption augmentation while enhancing the cross-modal retrieval ability with adaptive screening. As a fully data-driven process, sampling bias and empiricism that may confuse the cross-modal matching are avoided. We further encapsulate ExCae as a plug-in unit, which successfully improves the retrieval performance of current methods, validating ExCae’s flexibility and adaptability.
The contributions of this work are summarized as follows:
  • We propose eliminating the grunt workload of data acquisition and prompt design in caption augmentation through self-improvement and expertized selection, thereby mitigating empirical bias in supplemental texts while facilitating cross-modal matching without extra data.
  • We design an automatic cross-modal learning framework for video-text retrieval, in which video expressions are expanded using a Prompt engineer and are matched to texts under the guidance of learnable experts. The Prompt engineer facilitates generating effective and multi-angled captions while the learnable experts personalize the appropriate expression angles for cross-modal matching.
  • Experimental results show that our method outperforms (or is comparable to) current works even without using extra data, especially surpassing the best benchmark on the MSR-VTT by 3.8%. This paper also taps the application potentials of ExCae in video-text retrieval via extensive analytical and plug-in experiments.
Figure 2: Comparison of our method to the traditional
2 Related Work
Developing and pre-training video foundation models (VFMs) [11, 24] enhances the comprehension of video data, demonstrating efficacy in downstream tasks such as video-text retrieval. However, the effectiveness of VFMs heavily relies on huge amounts of data and computing resources. Much of the current work in video-text retrieval turns to a lighter modeling approach, adapting image-text retrieval models, such as CLIP [20], to video-text tasks. For instance, CLIP4Clip [15] and CLIP2TV [8] utilize the image knowledge encapsulated in CLIP, transforming it into effective video representations through different image aggregation networks. Besides, TS2-Net [14] dynamically adjusts the video token sequence, selecting tokens with richer information to enhance video representation. X-CLIP [16] employs an independent encoder for each frame, using a network that aggregates frames multiple times to generate the final video representation. DRL [23] introduces an additional loss to address the feature redundant issue. However, due to the information imbalance between videos and texts, these methods are stuck in insufficient cross-modal matching.
To reduce the cross-modal information gap, existing works, such as Cap4Video [26] and BIKE [27], enhance video-text matching via caption rewrite. They carefully design prompts, asking LLMs to rephrase the texts with empirically selective concepts, such as special objectives [18] or actions [17]. However, visual details remain deficient in the rewritten captions through these text-to-text literal restatements. More importantly, the hand-drafted lexicon and prompts severely limit the rewrite quality, motivating us to explore a self-learning caption enhancement method. To the best of our knowledge, this paper is a pioneer in investigating automatic caption enhancement to facilitate the comprehension of videos and texts.
3 Method
Unlike traditional text augmentation that rewrites captions on the text data side, our method (ExCae), automatically derives captions from videos by the CSI module and adaptively selects appropriate expressions (ECS module) to encode video for video-text matching (see Figure 2). Video and text embeddings would be aligned with standard cross-modal contrastive loss, as most existing SOTA methods [14, 15]. Both CSI and ECS are self-learning modules, whose details are introduced as follows.
3.1 Caption Self-improvement (CSI)
The objective of the CSI module is to learn the best prompt for caption generation. The main flow is illustrated in Figure 3, which is an iteration process gradually refining the query prompt. Specifically, a Captioner is asked to generate multiple captions to describe the video from different perspectives. We then evaluate the semantics of the generated captions, scoring them to update the prompt candidates’ pool. A Prompt engineer is introduced to rewrite the current best prompt, whose output is used as the latest query prompt of the Captioner. The prompt update iteration would be ended when no new candidates are added to the prompt candidates’ pool. (Pseudo code is provided in the Supplemental Materials.)
Caption scoring
Denote the query prompt of the Captioner as $p^{t}$, where $t$ records the iteration index. Correspondingly, the generated caption list for a video $v_{i}$ based on $p^{t}$ is written as $c^{t}_{i}$. As literal expressions, $c^{t}_{i}$ is evaluated from a semantic perspective. We advocate for the generation of captions that are semantically akin to the original texts, in order to mitigate the presence of cluttered and vacuous descriptions. However, merely preserving the semantic consistency may lead to generating a set of captions similar to the original text. This is opposite to the original intention of expression enhancement. To promote multi-angle descriptions, a diversity constraint that encourages variety within the caption list is supplemented. Suppose $K$ captions are contained in $c^{t}_{i}$, the caption generation score in the $t$ iteration round can be expressed as
$Score^{t}=\mathbb{E}_{i,k}[sim(\psi(c^{t}_{i,k}),\psi(t_{i}))]+div(\psi(c^{t}_{i}))$ (1)
where $t_{i}$ denotes the original text corresponding to $v_{i}$, $k\in[1,K]$ is the caption index in $c^{t}_{i}$. $\psi(\cdot)$ represents a semantic extractor, implemented by M3-embedding [5]. $sim(\cdot)$ and $div(\cdot)$ are similarity and diversity calculations (See Supplemental Materials for details). If $Score^{t}$ surpasses the historical records, $p^{t}$ would be added to the prompt candidates’ pool, and the best prompt is updated accordingly.
Prompt refinement
The current prompt engineering normally encodes the query prompts and adjusts the prompt tokens/embeddings through gradient descent [19, 32]. Although this parametric fine-tuning strategy is convenient for optimization, the adjusted embeddings could hardly be precisely re-projected to the text space, losing linguistic meaning and interpretability. Consequently, their generalization in video-text-related applications is limited. Inspired by language rewrites [7], this paper introduces a Prompt engineer to rewrite the query prompt. Take the current best prompt $p^{t}$ as input; the Prompt engineer is asked to rewrite $p^{t}$ with reference to the corresponding video-text pairs $(x,t)$ involved in this iteration. The rewritten prompt is then used as the query prompt for the Captioner in the next iteration, as shown in Figure 3.
Figure 3: Illustration of the CSI moudle
3.2 Expertized Caption Selection (ECS)
With the optimized CSI module, video-text pairs are updated as $((v,\tilde{c}),t)$, where $(v,\tilde{c})$ represents the video data $v$ and its generated captions $\tilde{c}$ using the optimal prompt, and $t$ denotes the original text corresponding to $v$. As displayed in Figure 2 (b), $(v,\tilde{c})$ and $t$ are the video-side and text-side input data, respectively. A straightforward approach for cross-modal representation learning is directly matching the encoded representations of $(v,\tilde{c})$ and $t$. However, the information dilation of video modality brings confounded knowledge while improving generalization, increasing the difficulty of cross-modal matching (see Figure 1).
Considering that the derived captions describe video content from multiple perspectives, we propose inserting an ECS module designed based on a mixture of experts to facilitate cross-modal matching. Write the outputs of visual encoders as $e_{v}=[e_{v1},e_{v2},...]$, which consists of $N+K$ elements, where $N$ is the number of sampled frames from videos, and $K$ is the number of derived captions. The ECS module is composed of multiple expert networks with learnable parameters ($[f_{1},f_{2},...,f_{M}]$), each of which refers to a description angle. $M$ denotes the total number of experts. A router $r_{m}$ is allocated to each expert $f_{m}$ as a gate, which is automatically learned to explore the most matched visual expression fusion for each video-text pair. Through this individualized video expression screening and fusion, the expression that is overly different from the corresponding texts is filtered on the video side, thereby narrowing the video-text modality gap. Mathematically, the visual codes are residually updated by a mixture of experts’ outputs:
$\tilde{e}_{v}=\frac{1}{M}\sum_{m}f_{m}(r_{m}(e_{v})\otimes e_{v})+e_{v}$ (2)
where $r_{m}(e_{v})$ outputs attention weights of $(N+K)$ video-side expressions for each sample. Specifically, a top-$R$ filter is applied on $r_{m}(e_{v})$, which activates partial expressions to learn a specific expert. $\otimes$ denotes the dot-product operation. As visual and textual embeddings will be aligned after further encoding by a shared encoder, parameters of the ECS module ($f_{m}$ and $r_{m}$) are optimized along with other encoding networks under the supervision of textual information.
Method Backbone Extra text-to-video video-to-text
[…]
Pre-trained foundation model
InternVideo [24] ViT-H/14 55.2 79.6 87.5 10.7 57.9 79.2 86.4 7.5
UMT-L [11] ViT-L/14 58.8 81.0 87.1 58.6 81.6 86.5
CLIP-VIP [31] ViT-B/16 57.7 80.5 88.2
mPLUG-2 [28] ViT-L/14 53.1 77.6 84.7
VAST [6] ViT-G/14 63.9 84.3 89.6
ViT-based
CLIP2TV [8] ViT-B/16 49.3 74.7 83.6 13.5 46.9 75.0 85.1 10.0
DRL [23] ViT-B/16 49.4 76.4 84.2 13.2 47.0 77.1 84.4 9.2
TS2-Net [14] ViT-B/16 47.8 76.8 85.2 13.7 47.8 76.0 84.6 8.5
Clip4Clip [15] ViT-B/16 46.4 72.1 82.0 14.7 45.4 73.4 82.4 10.7
X-CLIP [16] ViT-B/16 49.3 75.8 84.8 12.2 48.9 76.8 84.5 8.1
DMAE [9] ViT-B/16 49.9 75.8 85.5 12.5 49.6 76.3 85.0 8.5
Cap4Video [26] ViT-B/16 51.4 75.7 83.9 12.4 49.0 75.2 85.0 8.0
[…]
ViT-L/14 60.5 87.3 95.0 3.8 62.5 89.0 94.1 4.1
ViT-H/14 62.2 88.2 93.7 3.9 65.1 89.4 94.0 3.8
ViT-G/14 67.7 92.7 96.2 2.9 69.3 92.5 96.3 2.3
Table 1: Comparison with SOTA methods on MSR-VTT dataset.
4 Experiment
4.1 Datasets
MSR-VTT [29] is a large-scale dataset containing 10,000 video clips, each of which is annotated with 20 English sentences. Around 29K unique words are contained in all sentences. Following previous work [13], we use 9,000 clips for training and 1,000 clips for testing.
MSVD [3] consists of 1,970 videos with 80K captions. Each video contains around 40 English captions on average. The videos are split by 1,200/100/670 for training, validation and testing, respectively.
DiDeMo [1] contains over 10,000 videos collected from Flickr. Each video is described by natural language, with detailed information on camera movement, temporal transition indicators, and activities. The dataset is split into training, validation and test sets, each containing 8,395, 1,065 and 1,004 videos, respectively.
Figure 4: Examples of text-to-video retrieval results on MSR-VTT dataset. (The ground truths are marked in a red box.)
Method Backbone Extra text-to-video video-to-text
[…]
MSVD dataset
Pre-trained foundation model
InternVideo [24] ViT-H/14 58.4 84.5 90.4 8.2 76.3 96.8 98.7 2.1
UMT-L [11] ViT-L/14 58.2 83.9 89.6
ViT-based
CLIP2TV [8] ViT-B/16 50.2 79.8 87.9 8.6
DRL [23] ViT-B/16 50.0 81.5 89.5 68.7 92.5 95.6
Clip4Clip [15] ViT-B/16 47.2 77.7 9.1 63.2 87.2 4.2
X-CLIP [16] ViT-B/16 50.4 80.6 8.4 66.8 90.4 4.2
Cap4Video [26] ViT-B/16 51.8 80.8 88.3 8.3 70.0 93.2 96.2 2.4
[…]
ViT-L/14 53.6 80.3 86.9 10.4 67.6 90.9 96.5 3.1
ViT-H/14 56.5 82.2 88.4 9.4 74.1 93.4 97.2 2.3
ViT-G/14 59.2 83.4 90.7 5.1 76.9 96.6 99.0 1.6
DiDeMo dataset
Pre-trained foundation model
InternVideo [24] ViT-H/14 57.9 82.4 88.9 9.2 59.1 81.8 89.0 7.2
UMT-L [11] ViT-L/14 70.4 90.1 93.5 65.7 89.6 93.3
CLIP-VIP [31] ViT-B/16 55.3 82.0 89.3
mPLUG-2 [28] ViT-L/14 56.4 79.1 85.2
VAST [6] ViT-G/14 72.0 89.0 91.4
ViT-based
DRL [23] ViT-B/16 49.0 76.5 84.5 11.5 49.9 75.4 83.3 7.9
Clip4Clip [15] ViT-B/16 44.8 73.4 13.5 44.7 74.0 10.6
X-CLIP [16] ViT-B/16 47.8 79.3 12.6 47.8 76.8 10.5
[…]
ViT-L/14 56.9 80.9 87.9 6.5 58.8 83.2 90.0 4.8
ViT-H/14 60.1 83.3 90.0 5.9 61.0 85.6 92.3 4.5
ViT-G/14 62.0 85.5 92.3 4.9 63.7 87.9 93.3 4.1
Table 2: Comparison with SOTA methods on MSVD and DiDeMo datasets (underline marks the best ViT-based results)
4.2 Implementation Details
Preprocessing
Following the standard text preprocessing [20], each sentence in original texts and video-derived captions is split into word tokens with a max length of 70 using a CLIP tokenizer. For videos, we sample 8 frames from each item as visual data.
Model Setup
This paper reports the results of our method on multiple backbones, i.e., ViT-B/16, ViT-L/14, ViT-H/14, ViT-G/14, compared to SOTA video-text retrieval methods, showing the effect of model scale on performance as well. Ten captions are derived for each video. Sixteen experts are learned in the ECS module, two of which are activated to select appropriate representations in both training and inference processes. In the training phase, the CSI module is pre-learned to acquire the optimal prompt for caption generation, while the ECS module is co-trained alongside the backbone parameters. In the evaluation phase, multiple captions are derived from each video using the optimal prompt, together with the video itself as video-side expressions. (See Supplemental Materials for the learned best prompt and examples of video-derived captions.) The Captioner and Prompt engineer are implemented by GPT-4o, a large cross-modal inference model recently released by OpenAI in 2024. Besides the comparison experiment, we encapsulate ExCae as a plug-in unit and test its effect when introduced into existing video-text retrieval methods; specifically, we enhance video expression via CSI and professionalize the video-text matching process by ECS. Ablation studies and analytical experiments are conducted on MSR-VTT using ViT-G/14 as the base model.
Environment and Evaluation Metrics
Our model is implemented in PyTorch and trained on 8 NVIDIA A800 GPUs in a batch size of 16. We initialize the text and video encoder with pretrained weights from CLIP [20] and use the Adamax optimizer with a learning rate of 4e-6. Experimental results are evaluated by standard video-text retrieval metrics: Recall at Rank K (R@K, higher is better) and Mean Rank (MeanR, lower is better). R@K denotes the percentage of correct samples in the top K retrieved samples, where K is set to 1, 5 and 10, the same as in the previous works [31, 9]. MeanR records the average rank of correct items in the retrieved list.
4.3 Comparision to State-of-the-arts
Results of our method are compared to those of the SOTAs, including pre-trained foundation models like InternVideo [24], UMT-L [11], mPLUG-2 [28], VAST [6], and ViT-based methods that adapt the image-text models to the video-text task. Among the ViT-based methods, CLIP2TV [8], DRL [23], Clip4Clip [15] and CLIP-VIP [31] require the acquisition of auxiliary data for post-pretraining and/or adaptation. In contrast, TS-Net [14], X-CLIP [16], DMAE [9] and Cap4Video [26] are learned without extra data.
Results in Table 1 show that, on the MSR-VTT dataset, our method is superior not only to the ViT-based methods but also to the pre-training models relying on large-scale pre-training data. The results of existing methods show that the Top-1 recall accuracy could hardly transcend 50% without assistance from extra data. Instead, great improvement (up to 10%) is obtained when additional data is involved, either they are used for pre-training or fine-tuning. Despite not using any extra data, the Top-1 recall accuracy of our method (ExCae) boosts to 55.0% and 56.1% in text-to-video and video-to-text retrieval, respectively, when using ViT-B/16 as the base model. We are also encouraged by the observation that ExCae’s performance demonstrates an increase in correlation with the enlarged scale of the backbone model, resulting in an around 7% rise of R@1 from ViT-L/14 to Vit-G/14. Our results have comprehensively surpassed the current SOTA pre-training models, such as mPLUG-2 and VAST, when using the Vit-G/14 as the backbone. The rising trend of performance indicates the potential for further improvement of our method through scaling up the base model. Similar results are obtained in the experiments on MSVD and DiDeMo datasets, as recorded in Table 2. Our results are comparable to or even superior to those of pre-training models. Compared to existing ViT-based methods, ExCae shows a resounding victory, with Top-1 recall accuracy suppressing the best one by 7% on MSVD and 13% on DiDeMo. These results validate that our method is robust to cross-modal retrieval tasks for data collected from different fields.
We also demonstrate some video-text retrieval cases to present the performance of our method intuitively. The top three retrieved videos, when given a caption, are presented in Figure 4. In these cases, our method successfully picks the corresponding ground truth videos in the first place. The second and third hit videos are also highly related to the sentences in terms of content. Besides, Figure 5 compares our video-to-text retrieval results to the “Base” obtained without any caption enhancement. It is shown that the ground truth texts normally take the first rank when using our method. Compared to the base results, our method preferentially selects samples with semantics and descriptive perspectives similar to the ground truths. For example, our method successfully recognizes keywords of “voices”, “critic”, and “guitar” that are missed by the Base model. This is owing to the ECS module that adaptively learns the cross-modal matching for specific samples, benefiting the recognition of key semantics during video-text retrieval.
Figure 5: Examples of video-to-text retrieval results on MSR-VTT dataset. (The ground truths are marked in a red box.)
4.4 Plug-in Experiment
ExCae is encapsulated and supplemented with existing video-text retrieval methods, whose results are compared to those of the original. As shown in Table 3, our method successfully increases retrieval results of different video-text retrieval methods. The Top-1 recall accuracy is boosted by 4.8%, 9.4% and 7.7% for DRL, TSR-Net and Clip4Clip, respectively. It is observed that the results of plug-in experiments are similar to those obtained by our method with the same backbone (refer to Table 1). This indicates that ExCae has strong adaptability and expandability, which can flexibly combine with different models and constantly upgrade its performance as the technology evolves.
Method text-to-video video-to-text
R@1 R@5 R@10 R@1 R@5 R@10
DRL [23] 49.4 76.4 84.2 47.0 77.1 84.4
             + Ours 54.2 81.4 90.2 54.4 81.1 89.2
TS2-Net [14] 47.8 76.8 85.2 47.8 76.0 84.6
             + Ours 57.2 86.2 92.6 56.4 85.0 91.3
Clip4Clip [15] 46.4 72.1 82.0 45.4 73.4 82.4
             + Ours 54.1 82.1 89.5 51.2 80.6 88.8
Table 3: Results of plug-in experiments on MSR-VTT dataset.
Component text-to-video video-to-text
[…]
Video 49.9 78.1 88.1 7.9 49.6 78.8 87.3 7.3
52.9 78.5 86.7 10.1 51.3 78.5 87.1 10.3
51.9 79.7 88.3 9.1 51.4 79.8 87.1 10.9
Video-derived Caption 53.1 78.6 86.7 10.2 51.0 78.6 87.4 10.6
52.3 78.2 87.9 9.0 51.1 80.6 87.3 9.0
54.3 80.3 90.0 6.7 51.8 80.9 88.5 6.8
56.2 83.7 90.9 7.2 56.9 84.8 91.4 6.6
Video+ 63.2 90.1 95.2 3.3 65.0 89.9 94.6 3.3
Video-derived Caption 63.3 89.1 95.0 4.3 62.9 89.1 94.2 4.2
67.7 92.7 96.2 2.9 69.3 92.5 96.3 2.3
Table 4: Ablation studies of critical components
4.5 Ablation Studies
Our method expands the expression on the video side; thus, this section respectively evaluates the key components, CSI and ECS modules, under different input data cases. To test the effectiveness of the CSI module, we replace the optimal prompt with the initial one to simulate the ablation of the CSI module. The removal of the ECS module is realized by directly encoding all the video-side data as visual embeddings. As shown in Table 4, the R@1 is around 50% when only using the original video as video-side data without adding any components; this result is close to the existing ViT-based methods. The ECS module shows only slight positive effects on the original video. Similar results can be found if we merely change the video-side data from the original videos to video-derived captions. An obvious performance boost occurs when the original videos and video-derived captions are input in combination, validating the indispensability of these two data. In addition, we find that both CSI and ECS modules facilitate cross-modal retrieval. Despite in the absence of the other one, our method achieves over 6% gain in Top-1 recall accuracy. The effectiveness of the CSI and ECS modules proves the superiority of captions’ auto-enhancement and multi-perspective cross-modal matching. The best results are obtained by our full method, where both the CSI and ECS modules play mutually reinforcing roles in promoting video-text understanding.
4.6 Analytical Experiments
a) Effect of video-derived caption number ($N_{c}$)
Figure 6 plots the video-text retrieval results when raising $N_{c}$ from 1 to 10. The performance is really poor if only one caption is derived, with a Top-1 recall accuracy of lower than 30% on text-to-video retrieval. This result is inferior to any existing methods, reflecting that insufficient caption enhancement would even damage cross-modal learning. The reason is that all generated captions are forcibly activated to match the text if too few caption samples are input to the ECS module. Consequently, the original text may align with a description from a completely different perspective. Results are much better when increasing the number of generated captions to 3, achieving good performance close to the existing pre-training models. The performance further improves with the increase of the caption number, and the upward trend becomes stable when $N_{c}$ is greater than 7. This experiment suggests an approximate range of $[7,10]$ for the selection of video-derived caption numbers.
b) Effect of activated expert number ($N_{e}$)
Video-text retrieval results with different $N_{e}$ values are displayed in Figure 7. We also show the results obtained without the ECS module, labeled as 0 activated expert, as a reference. The retrieval results gradually improve as $N_{e}$ increases from 0 to 2. However, no further improvement in performance is achieved with the subsequent increase of activated experts. This may be because ECS would automatically favor partial experts and suppress the attention to less relevant expressions, although the majority of experts are activated.
Figure 6: Sensitivity to the number of video-derived captions
Figure 7: Sensitivity to the number of activated experts
5 Discussion
5.1 Superiority
ExCae attains remarkable video-text retrieval performance primarily due to its efficient narrowing of the video-text modality gap. As shown in Figure 8 (a), the representations of video-text pairs are drawn closer to each other through self-learning diffusional data augmentation (CSI) and self-adaptive cross-modal matching (ECS). The modality gap $||\Delta||$ [12] is reduced from 0.82 to 0.75, from 0.79 to 0.67, and from 0.84 to 0.74 on MSR-VTT, MSVD and DiDeMo, respectively. Moreover, ExCae has advantages in objectivity, generalization and self-adaptation. 1) ExCae is a data-driven method, in which the query prompts and caption selection are learned according to the training data. No additional human work (despite initialization), such as prompt design and lexicon establishment, is involved in the whole learning process. Thus, ExCae effectively avoids the adverse effects of empiricism. 2) As the Captioner is asked to generate multi-angle captions using a diversity constraint, data with more expression perspectives are provided for model learning. Therefore, ExCae inherits the strength of generalization improvement from data augmentation. 3) ExCae also has personalized ability owing to the parametric learning of the ECS module. The router in the ECS module would adaptively select the best expressions for cross-modal retrieval, thus maintaining the robustness of ExCae in different applications.
Figure 8: Modality gap and Convergence on MSR-VTT
5.2 Convergence
Since the generated captions are changeless once the optimal prompt is determined, the cross-modal post-training of our method owns the same convergence as existing ViT-based methods. Regarding learning the CSI module, the LLM would optimize query prompts in the given direction. With the increase of iteration, the semantic changes in query prompts become smaller, and thus the caption evaluation scores tend to be stable (See Figure 8 (b)). According to expertise from experiments, the CSI module tends to converge in 400 iterations.
5.3 Storage and Efficiency
Although a prompt optimization process is additionally involved, we believe that the storage and computation cost of ExCae is appreciable, especially compared to the pre-training methods. With regard to storage, no extra data or related lexicon needs to be stored. We only need to occupy a very small amount of storage space to record several generated captions in each epoch, which greatly reduces the storage cost. For the computation cost, our method is more efficient than the multi-modal foundation models without a pre-training process. In comparison to the ViT-based methods, the only additional computation is from the training of the CSI module. However, prompt optimization is a one-off training process, and we have also found that the learned prompt has generalization ability across different data sets. Therefore, the efficiency of our approach is comparable to existing video-text retrieval methods.
6 Conclusion
This paper proposes strengthening cross-modal understanding by enhancing video-text expressions in a self-learning manner, which avoids extra workload on data acquisition and prompt design. A CSI module that gradually refines query prompts is designed to generate high-quality captions with rich and diverse expressions, so as to expand the receptive field for cross-modal learning. In addition, multiple learnable experts are introduced as a filter to automatically select the appropriate perspectives of expressions for video-text matching. Our method is non-empirical and self-adaptive in comparison with current text rewriting methods. Experimental results demonstrate that our method achieves SOTA performance on benchmark datasets such as MSR-VTT, MSVD, and DiDemo. Looking ahead, we aspire to explore the variants adaptive to various cross-modal scenarios, thereby aiding a multitude of video-text-related tasks.

### v3

@@
 Abstract
-The burgeoning field of video-text retrieval has witnessed significant advancements with the advent of deep learning.
-However, the challenge of matching text and video persists due to inadequate textual descriptions of videos.
-The substantial information gap between the two modalities hinders a comprehensive understanding of videos, resulting in ambiguous retrieval results.
-While rewriting methods based on large language models have been proposed to broaden text expressions, carefully crafted prompts are essential to ensure the reasonableness and completeness of the rewritten texts.
-This paper proposes an automatic caption enhancement method that enhances expression quality and mitigates empiricism in augmented captions through self-learning.
-Additionally, an expertized caption selection mechanism is designed and introduced to customize augmented captions for each video, facilitating video-text matching.
+Video-text retrieval has been stuck in the information mismatch caused by personalized and inadequate textual descriptions of videos.
+The substantial information gap between the two modalities hinders an effective cross-modal representation alignment, resulting in ambiguous retrieval results.
+Although text rewriting methods have been proposed to broaden text expressions, the modality gap remains significant, as the text representation space is hardly expanded with insufficient semantic enrichment.
+Instead, this paper turns to enhancing visual presentation, bridging video expression closer to textual representation via caption generation and thereby facilitating video-text matching.
+While multimodal large language models (mLLM) have shown a powerful capability to convert video content into text, carefully crafted prompts are essential to ensure the reasonableness and completeness of the generated captions.
+Therefore, this paper proposes an automatic caption enhancement method that improves expression quality and mitigates empiricism in augmented captions through self-learning.
+Additionally, an expertized caption selection mechanism is designed and introduced to customize augmented captions for each video, further exploring the utilization potential of caption augmentation.
 Our method is entirely data-driven, which not only dispenses with heavy data collection and computation workload but also improves self-adaptability by circumventing lexicon dependence and introducing personalized matching.
 The superiority of our method is validated by state-of-the-art results on various benchmarks, specifically achieving Top-1 recall accuracy of 68.5% on MSR-VTT, 68.1% on MSVD, and 62.0% on DiDeMo.
-1 Introduction
+Our code is publicly available at https://github.com/CaryXiang/ECA4VTR.
+Index Terms:
+Video-Text Retrieval, Multimodal Representation Learning, Prompt Optimization.
+I Introduction
 Video-text understanding has become a new mainstream research with the population of short videos on social networks.
-In virtue of the learning ability of large models, multi-modal pre-training models [11, 6] are developed and achieve initial success in video-text retrieval tasks.
+In virtue of the learning ability of large models, multi-modal pre-training models [1, 2] are developed and achieve initial success in video-text retrieval (VTR) tasks.
 However, a stunning amount of data and computing resources are the essential basis to support the effectiveness of these pre-training models.
-Thus, it is argued that adapting the image-text pretrained model to video-text tasks is a lighter and more feasible approach [15, 31].
-They adapt the visual embeddings to the video field by incorporating temporal information [21, 4] or refining the visual-text matching strategies [8, 31].
-But unlike image-text pairs, video captions tend to be overly simplistic and one-sided, describing partial (even a signal) frames of the videos.
-These one-sided descriptions would be improperly matched to multifaceted video embeddings (Figure 1 (a)), resulting in biased cross-modal retrieval, and a bottleneck in performance improvement arises.
+A lighter and more feasible approach to achieve VTR is adapting the image-text pretrained model to video-text scenarios [3, 4].
+In this approach, the visual embeddings are adapted to the video field by introducing temporal information [5, 6] or refining the visual-text matching strategies [7, 4].
+But unlike image-text pairs, the textual descriptions of videos tend to be overly simplistic and one-sided, describing partial (even a single) frames of the videos.
+The improper alignment between one-sided descriptions and multifaceted video embeddings would confuse text understanding (Fig.
+1 (a)), resulting in biased cross-modal retrieval, and a bottleneck in performance improvement arises.
 Figure 1: Comprehension gap between videos and texts
-Frame sampling is one proper solution to handle the information imbalance issue.
+Frame sampling is one proper solution to handle the improper alignment issue.
 For instance, ClipBERT [8] sparsely samples video frames to meet the text descriptions.
-X-CLIP [16] further cuts text descriptions into words, employing cross-modal matching at both fine-grained (frame and word) and coarse-grained (video and sentence) levels.
+X-CLIP [9] further cuts text descriptions into words, searching for video frames for fine-grained cross-modal matching.
 As video information is sparsely divided and broken, these methods are limited in comprehensively utilizing and exploring deep relations across videos and texts.
-Another branch to reduce the information gap is data agumentation.
+Another branch to promote video-text understanding is data augmentation.
 Extra data, such as HD-VILA-100M [10] and LF-VILA [5], are collected to expand the receptive field for model adaptation, thereby enriching matching knowledge of the two modalities.
-To eliminate the heavy data collection workload, later research [27] rewrites captions by lexical analysis, in which attributes are extracted and emphasized to enhance expressive ability in text modality.
-Large language models (LLMs), such as Chat-GPT and LLaMA [22], are additionally introduced for the rewriting tasks [26, 7].
-Although up to 3% retrieval precision increase is achieved with the assistance of supplementary texts, the rewriting quality relies heavily on lexicon engineering [18] and hand-crafted prompts [17] (Figure 2 (a)).
-Recent works [2, 25] have been mired in the meticulous work of designing better prompts for caption augmentation, limited in human empiricism.
-This paper proposes an Expertized Caption auto-Enhancement (ExCae) method, which adaptively generates and personalized select captions for video-text retrieval, releasing extra work of data acquisition and avoiding the trouble of manual prompt design.
-Specifically, a Caption Self-improvement (CSI) module is designed to derive captions with rich content and diverse perspectives from videos.
-A Prompt engineer is introduced in the CSI module, which gradually adjusts and improves the query prompts via self-learning.
-With the enriched expressions on the video-side (video+video-derived captions), cross-modal matching becomes a more challenging task.
-Thus, we further develop an Expertized Caption Selection (ECS) module consisting of multiple learnable experts to automatically specify appropriate expressions for cross-modal matching.
-Seeing Figure 1 (b), video representation space dilates with caption augmentation and the cross-modal information gap is reduced after expertized selection.
+To address the substantial data collection burden, subsequent research has explored text rewriting strategies to enhance expressions on the text side.
+For example, Wu et al.
+[11] rewrite texts using lexical analysis, in which attributes are extracted to enhance expressive ability in the text modality.
+She et al.
+[12] further ameliorate text presentation using syntax-tree, emphasizing key words to assist better VTR.
+Besides generating textual descriptions, text enhancement is executed in the feature space [13, 14], providing embedding variants for cross-modal alignment.
+Despite achieving a retrieval precision increase of up to 3% with the aid of supplementary textual expressions, minimal additional semantics are introduced during the text-to-text rephrasing processes, which hinders further reduction of the modality gap.
+This restricts significant breakthroughs in cross-modal retrieval capability.
+Instead, this paper suggests enhancing the representations on the video side, deriving captions from videos to meet the expression modes and semantic information of texts.
+The rich semantics of videos are preserved, and video-text matching would be simplified due to the reduced modality gap achieved by this directional data augmentation.
+Although multimodal Large Language Models (mLLMs) have shown a decent ability in video understanding [15], generating appropriate video captions for VTR remains challenging, as it is revealed that the performance of mLLMs is query-sensitive.
+The writing quality of mLLMs relies heavily on lexicon engineering [16] and hand-crafted prompts [17].
+Recent studies [18, 19] have been mired in the meticulous work of designing better prompts for caption augmentation, limited in human empiricism.
+Inspired by automatic prompt optimization (APO) [20, 21], this paper proposes an Expertized Caption Auto-Enhancement (ExCae) method, as shown in Fig.
+2 (b).
+ExCae incorporates two main components: Caption Self-Improvement (CSI) and Expertized Caption Selection (ECS) modules, which adaptively generates and selectively personalizes captions for VTR.
+The two modules provide ExCae with the advantage of alleviating the additional burden of data acquisition and eliminating the complexities associated with manual prompt design.
+Specifically, in the CSI module, a Captioner implemented by a multimodal LLM (mLLM) is asked to derive captions from videos.
+We optimize the query prompt of the Captioner to ensure the generated captions contain rich content and diverse perspectives.
+To overcome the interpretability loss of prompt in existing gradient-based APO, this paper introduces a Prompt Engineer, which gradually improves the query prompts of the Captioner via repetitively inquest and output adjustment.
+With the enriched expressions on the video side (video+video-derived captions), the ECS module consisting of multiple learnable experts is designed to automatically specify appropriate expressions for cross-modal matching.
+Seeing Fig.
+1 (b), video representation space dilates with caption augmentation, and the cross-modal information gap is reduced after expertized selection.
 This indicates that the proposed method improves the generalization ability through caption augmentation while enhancing the cross-modal retrieval ability with adaptive screening.
@@
 The contributions of this work are summarized as follows:
-  • We propose eliminating the grunt workload of data acquisition and prompt design in caption augmentation through self-improvement and expertized selection, thereby mitigating empirical bias in supplemental texts while facilitating cross-modal matching without extra data.
-  • We design an automatic cross-modal learning framework for video-text retrieval, in which video expressions are expanded using a Prompt engineer and are matched to texts under the guidance of learnable experts.
-The Prompt engineer facilitates generating effective and multi-angled captions while the learnable experts personalize the appropriate expression angles for cross-modal matching.
+  • We propose reducing the modality comprehension gap for VTR by enhancing visual presentations via self-improved caption derivation from videos.
+Video captions are progressively refined, bridging video expression closer to text space while eliminating the grunt workload of data acquisition.
+  • We design an automatic video-text learning framework for VTR, in which video expressions are expanded using a Prompt Engineer and are matched with texts under the guidance of learnable experts.
+The Prompt Engineer facilitates effective and multi-angle caption generation through a self-improvement strategy that mitigates empirical bias in prompt design, while the learnable experts personalize the appropriate expression angles for cross-modal matching.
   • Experimental results show that our method outperforms (or is comparable to) current works even without using extra data, especially surpassing the best benchmark on the MSR-VTT by 3.8%.
@@
 Figure 2: Comparison of our method to the traditional
-2 Related Work
+II Related Work
 Developing and pre-training video foundation models (VFMs) [1, 22] enhances the comprehension of video data, demonstrating efficacy in downstream tasks such as video-text retrieval.
@@
 Much of the current work in video-text retrieval turns to a lighter modeling approach, adapting image-text retrieval models, such as CLIP [23], to video-text tasks.
-For instance, CLIP4Clip [15] and CLIP2TV [8] utilize the image knowledge encapsulated in CLIP, transforming it into effective video representations through different image aggregation networks.
-Besides, TS2-Net [14] dynamically adjusts the video token sequence, selecting tokens with richer information to enhance video representation.
-X-CLIP [16] employs an independent encoder for each frame, using a network that aggregates frames multiple times to generate the final video representation.
-DRL [23] introduces an additional loss to address the feature redundant issue.
-However, due to the information imbalance between videos and texts, these methods are stuck in insufficient cross-modal matching.
-To reduce the cross-modal information gap, existing works, such as Cap4Video [26] and BIKE [27], enhance video-text matching via caption rewrite.
-They carefully design prompts, asking LLMs to rephrase the texts with empirically selective concepts, such as special objectives [18] or actions [17].
-However, visual details remain deficient in the rewritten captions through these text-to-text literal restatements.
-More importantly, the hand-drafted lexicon and prompts severely limit the rewrite quality, motivating us to explore a self-learning caption enhancement method.
-To the best of our knowledge, this paper is a pioneer in investigating automatic caption enhancement to facilitate the comprehension of videos and texts.
-3 Method
-Unlike traditional text augmentation that rewrites captions on the text data side, our method (ExCae), automatically derives captions from videos by the CSI module and adaptively selects appropriate expressions (ECS module) to encode video for video-text matching (see Figure 2).
+For example, CLIP4Clip [3] and CLIP2TV [7] transform CLIP’s embeddings into effective video representations through various frame aggregation strategies.
+In addition, X-CLIP [9] proposes using an independent encoder for each frame to derive video representations.
+Video embeddings are further enhanced in TS2-Net [24], which dynamically selects video tokens with richer information.
+Transfer learning is another appropriate solution to adapt knowledge from image to video field, based on which Jin et al.
+[25] introduce a transformation block to adaptively transfer knowledge in the pre-trained CLIP from image-text to video-text.
+This adaptation is also achieved by a teacher-student framework in TeachClip [26].
+Considering the feature redundancy issue, DRL [27] designs a loss in channel decorrelation regularization to improve the quality of video representation.
+However, these methods suffer from ill-suited cross-modal matching due to the information imbalance between videos and texts.
+To improve cross-modal matching, existing works, such as BIKE [11] and SHE-NET [12], enhance text expression with lexical and syntax analysis.
+Although Large language models (LLMs), such as ChatGPT and LLaMA [28], are additionally introduced in the text rewriting tasks [29, 30, 31], visual details remain deficient in the rewritten texts through these text-to-text literal restatements.
+Generating captions from videos as auxiliaries is a brighter approach for text enhancement.
+These methods form new pairs using videos and derived captions to promote cross-modal representation learning [15].
+However, carefully designed prompts are necessary, and later works manually craft various query prompts, asking LLMs to describe empirically selective concepts, such as special objectives [16] or actions [17].
+The hand-drafted lexicon and prompts severely limit the rewrite quality, motivating us to explore a self-learning caption enhancement method.
+To the best of our knowledge, this paper is a pioneer in investigating automatic caption enhancement for videos to facilitate the comprehension of videos and texts.
+Figure 3: Illustration of the CSI module
+III Method
+Unlike traditional text augmentation that rewrites captions on the text side, our method (ExCae) automatically derives captions from videos by the CSI module and adaptively selects appropriate expressions (ECS module) to encode video for video-text matching (see Fig.
+2).
 Video and text embeddings would be aligned with standard cross-modal contrastive loss, as most existing SOTA methods [24, 3].
-Both CSI and ECS are self-learning modules, whose details are introduced as follows.
-3.1 Caption Self-improvement (CSI)
-The objective of the CSI module is to learn the best prompt for caption generation.
-The main flow is illustrated in Figure 3, which is an iteration process gradually refining the query prompt.
-Specifically, a Captioner is asked to generate multiple captions to describe the video from different perspectives.
-We then evaluate the semantics of the generated captions, scoring them to update the prompt candidates’ pool.
-A Prompt engineer is introduced to rewrite the current best prompt, whose output is used as the latest query prompt of the Captioner.
-The prompt update iteration would be ended when no new candidates are added to the prompt candidates’ pool.
-(Pseudo code is provided in the Supplemental Materials.)
+CSI and ECS are self-learning modules whose details are introduced as follows.
+III-A Caption Self-improvement (CSI)
+In the CSI module, a Captioner is included, which is an mLLM asked to generate a set of captions based on a query prompt to enhance video understanding.
+As intermediaries for video-text matching, the captions derived from videos are expected to be informative and text-oriented.
+To achieve this goal, CSI performs an iterative prompt refinement (Fig.
+3), gradually updating the query prompts of the Captioner by repeatedly inquiring a Prompt Engineer.
+In each iteration, the current query prompt is evaluated by caption scoring, an approximate assessment of VTR performance, resulting in the update of the prompt candidate.
+The Prompt engineer then rewrites the current best prompt, whose output is used as the latest query prompt of the Captioner.
+This prompt update iteration would end when no new candidates are added.
+Pseudo code is provided in Algorithm 1.
 Caption scoring
-Denote the query prompt of the Captioner as $p^{t}$, where $t$ records the iteration index.
-Correspondingly, the generated caption list for a video $v_{i}$ based on $p^{t}$ is written as $c^{t}_{i}$.
-As literal expressions, $c^{t}_{i}$ is evaluated from a semantic perspective.
-We advocate for the generation of captions that are semantically akin to the original texts, in order to mitigate the presence of cluttered and vacuous descriptions.
-However, merely preserving the semantic consistency may lead to generating a set of captions similar to the original text.
-This is opposite to the original intention of expression enhancement.
-To promote multi-angle descriptions, a diversity constraint that encourages variety within the caption list is supplemented.
-Suppose $K$ captions are contained in $c^{t}_{i}$, the caption generation score in the $t$ iteration round can be expressed as
-$Score^{t}=\mathbb{E}_{i,k}[sim(\psi(c^{t}_{i,k}),\psi(t_{i}))]+div(\psi(c^{t}_{i}))$ (1)
-where $t_{i}$ denotes the original text corresponding to $v_{i}$, $k\in[1,K]$ is the caption index in $c^{t}_{i}$.
-$\psi(\cdot)$ represents a semantic extractor, implemented by M3-embedding [5].
-$sim(\cdot)$ and $div(\cdot)$ are similarity and diversity calculations (See Supplemental Materials for details).
-If $Score^{t}$ surpasses the historical records, $p^{t}$ would be added to the prompt candidates’ pool, and the best prompt is updated accordingly.
+We hope that the adjustment of the query prompt will be positively correlated with the increase in VTR results throughout the iterative process.
+Directly evaluating VTR results as feedback for prompt refinement involves multimodal encoder optimization, where slight fluctuations in embeddings can significantly affect the direction of prompt optimization.
+Therefore, this paper designs a scoring mechanism that evaluates the quality of generated captions, implicitly reflecting the performance of VTR.
+Denote the query prompt for the Captioner at iteration $t$ as $p^{t}$.
+The generated caption list for video $v_{i}$, using $p^{t}$ is denoted as $\mathbf{c}_{i}^{t}=\{c_{i,1}^{t},c_{i,2}^{t},\ldots,c_{i,K}^{t}\}$, where $K$ is the number of captions per video.
+To balance semantic fidelity and diversity, $\mathbf{c}_{i}^{t}$ is evaluated from two perspectives:
+  1. 1.
+    Semantic Consistency: Ensures $\mathbf{c}_{i}^{t}$ semantically align with the original text $t_{i}$, mitigating the presence of cluttered and vacuous descriptions.
+  2. 2.
+    Caption Diversity: Encourages multi-angle descriptions in $\mathbf{c}_{i}^{t}$, avoiding redundancy to maintain the original intention of expression enhancement.
+The overall score for iteration $t$ is computed as:
+$Score^{t}=\mathbb{E}_{i,k}[sim(\psi(c^{t}_{i,k}),\psi(t_{i}))]+div(\psi(\mathbf{c}^{t}_{i}))$ (1)
+where $k\in[1,K]$ is the caption index, and $\psi(\cdot)$ represents a semantic extractor.
+$sim(\cdot)$ calculates the cosine similarity on semantics.
+$div(\cdot)$ quantifies diversity within $\mathbf{c}^{t}_{i})$:
+$div(\psi(\mathbf{c}^{t}_{i}))=\frac{1}{K(K-1)}\sum_{p\neq q}\left(1-sim\left(\psi(c_{i,p}^{t}),\psi(c_{i,q}^{t})\right)\right)$ (2)
+The higher value of $Score^{t}$ indicates better alignment with the original text and greater caption diversity.
+If $Score^{t}$ surpasses the historical records, the prompt candidate is updated by $p^{t}$, and the best prompt used for prompt refinement is updated accordingly.
 Prompt refinement
-The current prompt engineering normally encodes the query prompts and adjusts the prompt tokens/embeddings through gradient descent [19, 32].
+Current prompt engineering normally encodes the query prompts and adjusts the prompt tokens/embeddings through gradient descent [20, 21].
 Although this parametric fine-tuning strategy is convenient for optimization, the adjusted embeddings could hardly be precisely re-projected to the text space, losing linguistic meaning and interpretability.
@@
 Take the current best prompt $p^{t}$ as input; the Prompt engineer is asked to rewrite $p^{t}$ with reference to the corresponding video-text pairs $(x,t)$ involved in this iteration.
-The rewritten prompt is then used as the query prompt for the Captioner in the next iteration, as shown in Figure 3.
-Figure 3: Illustration of the CSI moudle
-3.2 Expertized Caption Selection (ECS)
+The rewritten prompt then inquires the Captioner in the next iteration, as shown in Fig.
+3.
+Algorithm 1 Caption Self-improvement (CSI) algorithm
+1:  Input: Video-text pairs $(x,t)$, Initial query prompt $p_{0}$, Maximum iteration $T$
+2:  Output: $Best\_prompt$
+3:  Initialization:
+4:      $Best\_prompt\leftarrow p_{0};$
+5:      $Best\_score\leftarrow-\infty;$
+6:  for $t=1,2,\dots,T$ do
+7:     $p_{t}\leftarrow PE(Best\_prompt)$;      // Refine prompt
+8:     if $Best\_prompt==p_{t}$ then
+9:        break;                           // No update
+10:     end if
+11:     $\mathbf{c}\leftarrow Captioner(x,p_{t})$;         // Generate Captions
+12:     Calculate $Score^{t}$ using Equation (1);
+13:     if $Score^{t}>Best\_score$ then
+14:        $Best\_score\leftarrow Score^{t};$
+15:        $Best\_prompt\leftarrow p_{t};$ ;
+16:     end if
+17:  end for
+III-B Expertized Caption Selection (ECS)
 With the optimized CSI module, video-text pairs are updated as $((v,\tilde{c}),t)$, where $(v,\tilde{c})$ represents the video data $v$ and its generated captions $\tilde{c}$ using the optimal prompt, and $t$ denotes the original text corresponding to $v$.
-As displayed in Figure 2 (b), $(v,\tilde{c})$ and $t$ are the video-side and text-side input data, respectively.
+As displayed in Fig.
+2 (b), $(v,\tilde{c})$ and $t$ are the video-side and text-side input data, respectively.
 A straightforward approach for cross-modal representation learning is directly matching the encoded representations of $(v,\tilde{c})$ and $t$.
-However, the information dilation of video modality brings confounded knowledge while improving generalization, increasing the difficulty of cross-modal matching (see Figure 1).
+However, the information dilation of video modality brings confounded knowledge while improving generalization, increasing the difficulty of cross-modal matching (see Fig.
+1).
 Considering that the derived captions describe video content from multiple perspectives, we propose inserting an ECS module designed based on a mixture of experts to facilitate cross-modal matching.
@@
 As visual and textual embeddings will be aligned after further encoding by a shared encoder, parameters of the ECS module ($f_{m}$ and $r_{m}$) are optimized along with other encoding networks under the supervision of textual information.
+TABLE I: Comparison with SOTA methods on MSR-VTT dataset (underline marks the best ViT-based results)
 Method Backbone Extra text-to-video video-to-text
@@
 DMAE [33] ViT-B/16 49.9 75.8 85.5 12.5 49.6 76.3 85.0 8.5
-Cap4Video [26] ViT-B/16 51.4 75.7 83.9 12.4 49.0 75.2 85.0 8.0
+Cap4Video++ [15] ViT-B/16 52.3 76.8 85.8 11.5 50.0 75.9 86.0 7.8
+TeachClip [26] ViT-B/16 48.0 75.9 83.5
 […]
@@
 ViT-G/14 67.7 92.7 96.2 2.9 69.3 92.5 96.3 2.3
-Table 1: Comparison with SOTA methods on MSR-VTT dataset.
-4 Experiment
-4.1 Datasets
-MSR-VTT [29] is a large-scale dataset containing 10,000 video clips, each of which is annotated with 20 English sentences.
-Around 29K unique words are contained in all sentences.
-Following previous work [13], we use 9,000 clips for training and 1,000 clips for testing.
-MSVD [3] consists of 1,970 videos with 80K captions.
-Each video contains around 40 English captions on average.
-The videos are split by 1,200/100/670 for training, validation and testing, respectively.
-DiDeMo [1] contains over 10,000 videos collected from Flickr.
-Each video is described by natural language, with detailed information on camera movement, temporal transition indicators, and activities.
-The dataset is split into training, validation and test sets, each containing 8,395, 1,065 and 1,004 videos, respectively.
-Figure 4: Examples of text-to-video retrieval results on MSR-VTT dataset. (The ground truths are marked in a red box.)
+IV Experiment
+IV-A Datasets
+MSR-VTT [34] is a large-scale dataset containing 10,000 video clips.
+Each video is described in 20 English sentences.
+This paper uses 9,000 clips for training and 1,000 clips for testing, as in the setup of previous work [35].
+MSVD [36] consists of 1,970 videos.
+Each video contains around 40 English captions.
+In total, around 80K captions are provided.
+The dataset is derived by 1,200/100/670 for training, validation and testing, respectively.
+DiDeMo [37] contains over 10,000 videos downloaded from Flickr.
+Each video has a caption recording detailed information on camera movement, temporal transition indicators, and activities.
+We split the dataset into training, validation and test sets, each containing 8,395, 1,065 and 1,004 videos, respectively.
+TABLE II: Comparison with SOTA methods on MSVD and DiDeMo datasets (underline marks the best ViT-based results)
 Method Backbone Extra text-to-video video-to-text
@@
 X-CLIP [9] ViT-B/16 50.4 80.6 8.4 66.8 90.4 4.2
-Cap4Video [26] ViT-B/16 51.8 80.8 88.3 8.3 70.0 93.2 96.2 2.4
+Cap4Video++ [15] ViT-B/16 51.8 80.8 88.3 8.3 70.0 93.2 96.2 2.4
 […]
@@
 X-CLIP [9] ViT-B/16 47.8 79.3 12.6 47.8 76.8 10.5
+Cap4Video++ [15] ViT-B/16 52.5 80.0 87.0 10.3 51.2 78.5 87.4 7.3
+TeachClip [26] ViT-B/16 43.7 72.7
 […]
@@
 ViT-G/14 62.0 85.5 92.3 4.9 63.7 87.9 93.3 4.1
-Table 2: Comparison with SOTA methods on MSVD and DiDeMo datasets (underline marks the best ViT-based results)
-4.2 Implementation Details
+IV-B Implementation Details
 Preprocessing
-Following the standard text preprocessing [20], each sentence in original texts and video-derived captions is split into word tokens with a max length of 70 using a CLIP tokenizer.
+Following the standard text preprocessing [23], we split the textual data, including original texts and video-derived captions, into word tokens with a max length of 70 using a CLIP tokenizer.
 For videos, we sample 8 frames from each item as visual data.
@@
 In the evaluation phase, multiple captions are derived from each video using the optimal prompt, together with the video itself as video-side expressions.
-(See Supplemental Materials for the learned best prompt and examples of video-derived captions.) The Captioner and Prompt engineer are implemented by GPT-4o, a large cross-modal inference model recently released by OpenAI in 2024.
-Besides the comparison experiment, we encapsulate ExCae as a plug-in unit and test its effect when introduced into existing video-text retrieval methods; specifically, we enhance video expression via CSI and professionalize the video-text matching process by ECS.
+The Captioner and the Prompt engineer are implemented by GPT-4o, a large cross-modal inference model recently released by OpenAI in 2024.
+The initial prompt asking the Captioner is set as “Generate 10 captions from different perspectives about this video.” The semantic embedding extractor is Equation (1) is implemented by M3-embedding [38].
+In addition to the comparison experiment, we encapsulate ExCae as a plug-in unit and test its effect when introduced into existing video-text retrieval methods; specifically, we enhance video expression via CSI and professionalize the video-text matching process by ECS.
 Ablation studies and analytical experiments are conducted on MSR-VTT using ViT-G/14 as the base model.
 Environment and Evaluation Metrics
-Our model is implemented in PyTorch and trained on 8 NVIDIA A800 GPUs in a batch size of 16.
-We initialize the text and video encoder with pretrained weights from CLIP [20] and use the Adamax optimizer with a learning rate of 4e-6.
-Experimental results are evaluated by standard video-text retrieval metrics: Recall at Rank K (R@K, higher is better) and Mean Rank (MeanR, lower is better).
-R@K denotes the percentage of correct samples in the top K retrieved samples, where K is set to 1, 5 and 10, the same as in the previous works [31, 9].
+Our model is trained on 8 NVIDIA A800 GPUs in a batch size of 16.
+The pretrained weights from CLIP [23] are used to initialize the text and video encoder in ExCae.
+The model is trained using the Adamax optimizer with a learning rate of 4e-6.
+Experimental results are evaluated by standard video-text retrieval metrics: Recall at Rank K (R@K) and Mean Rank (MeanR).
+R@K denotes the percentage of correct samples in the top K retrieved samples.
+In this paper, K is set to 1, 5 and 10, respectively, the same as in previous works [4, 33].
 MeanR records the average rank of correct items in the retrieved list.
-4.3 Comparision to State-of-the-arts
-Results of our method are compared to those of the SOTAs, including pre-trained foundation models like InternVideo [24], UMT-L [11], mPLUG-2 [28], VAST [6], and ViT-based methods that adapt the image-text models to the video-text task.
+The higher value in R@K means better retrieval performance.
+In contrast, a lower MeanR is preferred in VTR.
+IV-C Comparision to State-of-the-arts
+The results of our method are compared to those of the SOTAs, including pre-trained foundation models like InternVideo [22], UMT-L [1], mPLUG-2 [32], VAST [2], and ViT-based methods that adapt the image-text models to the video-text task.
 Among the ViT-based methods, CLIP2TV [7], DRL [27], Clip4Clip [3] and CLIP-VIP [4] require the acquisition of auxiliary data for post-pretraining and/or adaptation.
-In contrast, TS-Net [14], X-CLIP [16], DMAE [9] and Cap4Video [26] are learned without extra data.
-Results in Table 1 show that, on the MSR-VTT dataset, our method is superior not only to the ViT-based methods but also to the pre-training models relying on large-scale pre-training data.
+In contrast, TS-Net [24], X-CLIP [9], DMAE [33], Cap4Video++ [15] and TeachClip [26] are learned without extra data.
+Results in Table I show that, on the MSR-VTT dataset, our method is superior to all existing methods, including pre-training foundation models and the ViT-based methods.
 The results of existing methods show that the Top-1 recall accuracy could hardly transcend 50% without assistance from extra data.
@@
 The rising trend of performance indicates the potential for further improvement of our method through scaling up the base model.
-Similar results are obtained in the experiments on MSVD and DiDeMo datasets, as recorded in Table 2.
+Similar results are obtained in experiments on MSVD and DiDeMo datasets, as recorded in Table II.
 Our results are comparable to or even superior to those of pre-training models.
 Compared to existing ViT-based methods, ExCae shows a resounding victory, with Top-1 recall accuracy suppressing the best one by 7% on MSVD and 13% on DiDeMo.
-These results validate that our method is robust to cross-modal retrieval tasks for data collected from different fields.
-We also demonstrate some video-text retrieval cases to present the performance of our method intuitively.
-The top three retrieved videos, when given a caption, are presented in Figure 4.
-In these cases, our method successfully picks the corresponding ground truth videos in the first place.
-The second and third hit videos are also highly related to the sentences in terms of content.
-Besides, Figure 5 compares our video-to-text retrieval results to the “Base” obtained without any caption enhancement.
-It is shown that the ground truth texts normally take the first rank when using our method.
+These results validate that ExCae is robust to VRT tasks for different data sources.
+Moreover, we demonstrate several VTR cases to show our results more intuitively.
+Fig.
+4 presents the top three videos retrieved when given a specific caption.
+It is shown that our method successfully picks the ground-truth videos in the first place.
+We appreciate seeing that the second and third hit videos are also highly related to the sentences in terms of content.
+In Fig.
+5, we display the text located at the top of the retrieval list when providing a video.
+Our results are compared to the “Base” obtained without any caption enhancement.
+The ground-truth texts normally take the first rank when using our method.
 Compared to the base results, our method preferentially selects samples with semantics and descriptive perspectives similar to the ground truths.
@@
 This is owing to the ECS module that adaptively learns the cross-modal matching for specific samples, benefiting the recognition of key semantics during video-text retrieval.
+Figure 4: Examples of text-to-video retrieval results on MSR-VTT dataset. (The ground truths are marked in a red box.)
 Figure 5: Examples of video-to-text retrieval results on MSR-VTT dataset. (The ground truths are marked in a red box.)
-4.4 Plug-in Experiment
+IV-D Plug-in Experiment
 ExCae is encapsulated and supplemented with existing video-text retrieval methods, whose results are compared to those of the original.
-As shown in Table 3, our method successfully increases retrieval results of different video-text retrieval methods.
+As shown in Table III, our method successfully increases retrieval results of different video-text retrieval methods.
 The Top-1 recall accuracy is boosted by 4.8%, 9.4% and 7.7% for DRL, TSR-Net and Clip4Clip, respectively.
-It is observed that the results of plug-in experiments are similar to those obtained by our method with the same backbone (refer to Table 1).
+It is observed that the results of plug-in experiments are similar to those obtained by our method with the same backbone (refer to Table I).
 This indicates that ExCae has strong adaptability and expandability, which can flexibly combine with different models and constantly upgrade its performance as the technology evolves.
+IV-E Ablation Studies
+Our method expands the expression on the video side; thus, this section respectively evaluates the key components, CSI and ECS modules, under different input data cases.
+To test the effectiveness of the CSI module, we replace the optimal prompt with the initial one to simulate the ablation of the CSI module.
+The removal of the ECS module is realized by directly encoding all the video-side data as visual embeddings.
+As shown in Table IV, the R@1 is around 50% when only using the original video as video-side data without adding any components; this result is close to the existing ViT-based methods.
+The ECS module shows only slight positive effects on the original video.
+Similar results can be found if we merely change the video-side data from the original videos to video-derived captions.
+An obvious performance boost occurs when the original videos and video-derived captions are input in combination, validating the indispensability of these two data.
+In addition, we find that both CSI and ECS modules facilitate cross-modal retrieval.
+Despite in the absence of the other one, our method achieves over 6% gain in Top-1 recall accuracy.
+The effectiveness of the CSI and ECS modules proves the superiority of captions’ auto-enhancement and multi-perspective cross-modal matching.
+The best results are obtained by our full method, where both the CSI and ECS modules play mutually reinforcing roles in promoting video-text understanding.
+IV-F Analytical Experiments
+TABLE III: Results of plug-in experiments on MSR-VTT dataset.
 Method text-to-video video-to-text
@@
              + Ours 54.1 82.1 89.5 51.2 80.6 88.8
-Table 3: Results of plug-in experiments on MSR-VTT dataset.
+Effect of video-derived caption number ($N_{c}$)
+Fig.
+6 plots the video-text retrieval results when raising $N_{c}$ from 1 to 10.
+The performance is really poor if only one caption is derived, with a Top-1 recall accuracy of lower than 30% on text-to-video retrieval.
+This result is inferior to any existing methods, reflecting that insufficient caption enhancement would even damage cross-modal learning.
+The reason is that all generated captions are forcibly activated to match the text if too few caption samples are input to the ECS module.
+Consequently, the original text may align with a description from a completely different perspective.
+Results are much better when increasing the number of generated captions to 3, achieving good performance close to the existing pre-training models.
+The performance further improves with the increase of the caption number, and the upward trend becomes stable when $N_{c}$ is greater than 7.
+This experiment suggests an approximate range of $[7,10]$ for the selection of video-derived caption numbers.
+Effect of activated expert number ($N_{e}$)
+Video-text retrieval results with different $N_{e}$ values are displayed in Fig.
+7.
+We also show the results obtained without the ECS module, labeled as 0 activated expert, as a reference.
+The retrieval results gradually improve as $N_{e}$ increases from 0 to 2.
+However, no further improvement in performance is achieved with the subsequent increase of activated experts.
+This may be because ECS would automatically favor partial experts and suppress the attention to less relevant expressions, although the majority of experts are activated.
+TABLE IV: Ablation studies of critical components
 Component text-to-video video-to-text
@@
 67.7 92.7 96.2 2.9 69.3 92.5 96.3 2.3
-Table 4: Ablation studies of critical components
-4.5 Ablation Studies
-Our method expands the expression on the video side; thus, this section respectively evaluates the key components, CSI and ECS modules, under different input data cases.
-To test the effectiveness of the CSI module, we replace the optimal prompt with the initial one to simulate the ablation of the CSI module.
-The removal of the ECS module is realized by directly encoding all the video-side data as visual embeddings.
-As shown in Table 4, the R@1 is around 50% when only using the original video as video-side data without adding any components; this result is close to the existing ViT-based methods.
-The ECS module shows only slight positive effects on the original video.
-Similar results can be found if we merely change the video-side data from the original videos to video-derived captions.
-An obvious performance boost occurs when the original videos and video-derived captions are input in combination, validating the indispensability of these two data.
-In addition, we find that both CSI and ECS modules facilitate cross-modal retrieval.
-Despite in the absence of the other one, our method achieves over 6% gain in Top-1 recall accuracy.
-The effectiveness of the CSI and ECS modules proves the superiority of captions’ auto-enhancement and multi-perspective cross-modal matching.
-The best results are obtained by our full method, where both the CSI and ECS modules play mutually reinforcing roles in promoting video-text understanding.
-4.6 Analytical Experiments
-a) Effect of video-derived caption number ($N_{c}$)
-Figure 6 plots the video-text retrieval results when raising $N_{c}$ from 1 to 10.
-The performance is really poor if only one caption is derived, with a Top-1 recall accuracy of lower than 30% on text-to-video retrieval.
-This result is inferior to any existing methods, reflecting that insufficient caption enhancement would even damage cross-modal learning.
-The reason is that all generated captions are forcibly activated to match the text if too few caption samples are input to the ECS module.
-Consequently, the original text may align with a description from a completely different perspective.
-Results are much better when increasing the number of generated captions to 3, achieving good performance close to the existing pre-training models.
-The performance further improves with the increase of the caption number, and the upward trend becomes stable when $N_{c}$ is greater than 7.
-This experiment suggests an approximate range of $[7,10]$ for the selection of video-derived caption numbers.
-b) Effect of activated expert number ($N_{e}$)
-Video-text retrieval results with different $N_{e}$ values are displayed in Figure 7.
-We also show the results obtained without the ECS module, labeled as 0 activated expert, as a reference.
-The retrieval results gradually improve as $N_{e}$ increases from 0 to 2.
-However, no further improvement in performance is achieved with the subsequent increase of activated experts.
-This may be because ECS would automatically favor partial experts and suppress the attention to less relevant expressions, although the majority of experts are activated.
 Figure 6: Sensitivity to the number of video-derived captions
 Figure 7: Sensitivity to the number of activated experts
-5 Discussion
-5.1 Superiority
+V Discussion
+V-A Superiority
 ExCae attains remarkable video-text retrieval performance primarily due to its efficient narrowing of the video-text modality gap.
-As shown in Figure 8 (a), the representations of video-text pairs are drawn closer to each other through self-learning diffusional data augmentation (CSI) and self-adaptive cross-modal matching (ECS).
+As shown in Fig.
+8 (a), the representations of video-text pairs are drawn closer to each other through self-learning diffusional data augmentation (CSI) and self-adaptive cross-modal matching (ECS).
 The modality gap $||\Delta||$ [39] is reduced from 0.82 to 0.75, from 0.79 to 0.67, and from 0.84 to 0.74 on MSR-VTT, MSVD and DiDeMo, respectively.
@@
 Figure 8: Modality gap and Convergence on MSR-VTT
-5.2 Convergence
+V-B Convergence
 Since the generated captions are changeless once the optimal prompt is determined, the cross-modal post-training of our method owns the same convergence as existing ViT-based methods.
 Regarding learning the CSI module, the LLM would optimize query prompts in the given direction.
-With the increase of iteration, the semantic changes in query prompts become smaller, and thus the caption evaluation scores tend to be stable (See Figure 8 (b)).
+With the increase of iteration, the semantic changes in query prompts become smaller, and thus the caption evaluation scores tend to be stable (See Fig.
+8 (b)).
 According to expertise from experiments, the CSI module tends to converge in 400 iterations.
-5.3 Storage and Efficiency
+V-C Storage and Efficiency
 Although a prompt optimization process is additionally involved, we believe that the storage and computation cost of ExCae is appreciable, especially compared to the pre-training methods.
@@
 However, prompt optimization is a one-off training process, and we have also found that the learned prompt has generalization ability across different data sets.
-Therefore, the efficiency of our approach is comparable to existing video-text retrieval methods.
-6 Conclusion
-This paper proposes strengthening cross-modal understanding by enhancing video-text expressions in a self-learning manner, which avoids extra workload on data acquisition and prompt design.
-A CSI module that gradually refines query prompts is designed to generate high-quality captions with rich and diverse expressions, so as to expand the receptive field for cross-modal learning.
+Therefore, the efficiency of our approach is comparable to existing VTR methods.
+VI Conclusion
+This paper proposes strengthening cross-modal understanding by enhancing video expressions rather than text rewriting, which effectively reduces the modality gap without additional workload on data acquisition.
+We innovatively developed a CSI module that automatically refines query prompts to generate high-quality video captions with rich and diverse expressions.
+The visual domain is expanded to meet the expression modes and semantic information of texts, so as to facilitate cross-modal understanding.
 In addition, multiple learnable experts are introduced as a filter to automatically select the appropriate perspectives of expressions for video-text matching.
@@
 Looking ahead, we aspire to explore the variants adaptive to various cross-modal scenarios, thereby aiding a multitude of video-text-related tasks.
+Acknowledgments
+This work was supported in part by the National Natural Science Foundation of China (NSFC) under Grant 62472105, in part by the Natural Science Foundation of Guangdong Province under Grants 2024A1515010186 and 2025A1515011385, as well as the Regional Joint Fund Project of the Basic and Applied Basic Research Foundation of Guangdong Province under Grant 2022A1515140096.